# On completeness of Foliated structures, and Null Killing fields

### v1

@@ 1.3.2. Affine lightlike geometry
One could relax the $({\sf{L}}_{\sf{u}}(n),\mathbb{R}^{n+1})$-geometry by assuming instead of an invariant vector field, only an invariant line field. This geometry also comes from concrete examples known as “weakly pp-waves” (see Example 3.1). In this case, the structure group is
@@
We call this $({\sf{L}}(n),\mathbb{R}^{n+1})$-geometry the affine lightlike geometry.
This geometry, however, doesn’t satisfy that any compact $({\sf{L}}(n),\mathbb{R}^{n+1})$-manifold is complete. An incomplete example can be obtained simply by taking the product of a Hopf circle and a Euclidean torus. In this case, and using essentially the same techniques as for Theorem 1.9, we characterize the completeness in the following way
@@
Let $(M,\mathfrak{F})$ be a $(\mathsf{L}(n),\mathbb{R}^{n+1})$-foliated compact manifold. The leaves of $\mathfrak{F}$ are $(\mathsf{L}(n),\mathbb{R}^{n+1})$-complete if and only if all geodesics tangent to the $\mathcal{V}$-leaves are affinely (or geodesically) complete , where $\mathcal{V}$ is the $1$-dimensional foliation tangent to the line field induced on any $\mathfrak{F}$-leaf by this structure.
We give an explicit example showing that there are $(\mathsf{L}(n),\mathbb{R}^{n+1})$-foliated compact manifolds, where the leaves are not complete.
@@
A natural generalization of the above situation is the following pseudo-Riemannian variant of the $({\sf{L}}_{\sf{u}}(n,0),\mathbb{R}^{n+1})$-geometry. Namely,
@@
where $\mathsf{O}(p,q)$ is the orthogonal group of the pseudo-Riemannian scalar product of signature $(p,q)$. The completeness of a compact $({\sf{L}}_{\sf{u}}(p,q),\mathbb{R}^{p+q+1})$-manifold (and its dual geometry) is itself an open question. In a recent paper [3], the author shows that the answer is positive for the $({\sf{L}}_{\sf{u}}(n-1,1),\mathbb{R}^{n+1})$-geometry, i.e. a compact $({\sf{L}}_{\sf{u}}(n-1,1),\mathbb{R}^{n+1})$-manifold is complete. But it turns out that the foliated completeness Problem 1.8 fails for this geometry: we give an example (Example 4.4) of a compact $({\sf{L}}_{\sf{u}}(n-1,1),\mathbb{R}^{n+1})$-foliated manifold with incomplete leaves, for any $n\geq 2$.
For the dual geometry we show
@@
Let $M$ be a compact $({\sf{L}}_{\sf{u}}^{*}(n-1,1),\mathbb{R}^{n+1})$-manifold. If $\pi_{1}(M)$ is virtually solvable, then $M$ is complete.
From the incomplete Example 4.4, one can in fact obtain compact $({\sf{L}}_{\sf{u}}(p,q),\mathbb{R}^{p+q+1})$-foliated manifolds as well as compact $({\sf{L}}_{\sf{u}}^{*}(p,q),\mathbb{R}^{p+q+1})$-manifolds with incomplete leaves, when $p,q\geq 1$.
Remark 1.13.
Note that for the compact $({\sf{L}}_{\sf{u}}(n-1,1),\mathbb{R}^{n+1})$-foliated manifolds ($n\geq 2$) with incomplete leaves, constructed in Example 4.4, the structure group preserves a Lorentzian metric. Thus, the question is still open when the holonomy doesn’t preserve a Lorentzian metric. The Latter case naturally arises if we consider “generalized” pp-waves. Namely, they are manifolds with a metric of index $(n,2)$, which have a parallel vector field $V$ with a flat $V^{\perp}$-distribution.
@@
The organization of the paper is different from the order presented in the introduction. Section 2 deals with the $(G,X)$-foliated completeness problem in Item (B), in the case of the affine unimodular lightlike geometry and its dual. We prove Theorem 1.9 and Theorem 1.10 which give leafwise completeness results. Then we study the case of the affine lightlike geometry in Section 3: we prove Theorem 1.11 that characterizes the leafwise completeness, and give leafwise-incomplete examples. Section 4 deals with the pseudo-Riemannian variants of the previous structures. In Section 5, we deal with the foliated completeness problem in Item (A), and prove Theorem 5.1. Finally, Section 6 is devoted to the study of the global completeness question in the presence of a null Killing field, focusing on dimension $3$ in Section 7.
Acknowledgement
We would like to warmly thank Abdelghani Zeghib and Souheib Allout for the helpful discussions, that made this work possible. The first author is also grateful to Ines Kath for her support and encouragements, and for the helpful remarks on the early drafts of this paper. The second author would like to warmly thank Miguel Sanchez for his interest in her work, and for pointing out to her that there is no known incomplete compact Lorentzian manifold with a null Killing field, which motivated this part of the work.
The second author is partially supported by the grants, PID2020-116126GB-I00
@@ 2. $(G,X)$-foliated completeness problem: affine unimodular lightlike geometry
Let $G$ be a Lie group, and $X$ a space on which $G$ acts smoothly transitively, then, one has $X=G/H$, where $H$ is a closed subgroup of $G$. A $(G,X)$-structure on $M$ is an atlas on $M$, where the chart transitions are restrictions of elements of $G$. A manifold with a $(G,X)$-structure is called a $(G,X)$-manifold. A $(G,X)$-manifold $M$ gives rise to a local diffeomorphism $D$, called the developing map, from the universal cover to the model space $X$, together with a representation $\rho:\pi_{1}(M)\to G$, called the holonomy representation (see [10, p. 100] or [23, p. 34]). Moreover, the developing map and the holonomy representation satisfy the following equivariance property
$D(\gamma x)=\rho(\gamma)D(x),$
@@
Let $M$ be a $(G,X)$-manifold, we say that $M$ is complete if the developing map $D:\widetilde{M}\to X$ is a covering map.
@@ 2.2. $(G,X)$-foliation
We introduce here the notion of a $(G,X)$-foliation. Examples are given in the next sections. A smooth manifold $M$ with a foliation $\mathfrak{F}$ will be called a foliated manifold and denoted by a couple $(M,\mathfrak{F})$.
Definition 2.3 ($(G,X)$-foliation).
@@
We say that a manifold $M$ has a $(G,X)$-foliation if $M$ is a foliated manifold such that the leaves of $\mathfrak{F}$ are modeled on the $(G,X)$-geometry. More precisely, for every $p\in M$ there exist an open neighborhood $U$ and a diffeomorphism (into the image) $\varphi:U\to I\times X$ such that every leaf the foliation of $U$ (induced by $\mathfrak{F}$) is mapped to a level $(t,V_{t})\subset I\times X$. And if $(\varphi_{i},U_{i})$ and $(\varphi_{j},U_{j})$ are two such maps, the transition map $\varphi_{ij}:=\varphi_{j}\varphi_{i}^{-1}$ on each level $(t,V_{t})\subset\varphi_{i}(U_{i}\cap U_{j})$ is the restriction of an element $g_{t}\in G$. Moreover, the $g_{t}$ (necessarily unique) varies continuously from a level to another.
@@
Pullback of $G$-invariant tensor fields. Consider $(M,\mathfrak{F})$ a $(G,X)$-foliated manifold. Let $T$ be a tensor field on $X$ which is $G$-invariant. One can associate a tensor field $T_{M}$ on $M$ in the following way. For any point $p\in M$, we take a chart around $p$ and we pullback the tensor field $T$ in a neighborhood of $p$. We claim that we get a well defined tensor field $T_{M}$ on $M$. Indeed, the transition maps, when restricted to a leaf, are restrictions of an element of $G$, and $T$ is $G$-invariant. Moreover, by construction, $T_{M}$ is continuous on $M$. We will refer to $T_{M}$ as the “pullback” of $T$.
@@
Recall that the affine unimodular lightlike geometry is the one modeled on $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$, where $\mathsf{L}_{\sf{u}}(n)$ is the subgroup of $\sf{Aff}(\mathbb{R}^{n+1})$ preserving the degenerate Riemannian metric $h:=dx_{2}^{2}+dx_{3}^{2}+..+dx_{n+1}^{2}$, and the isotropic vector field $V:=\partial_{x_{1}}$. In this subsection, $V$ and $h$ will always refer to these ${\sf{L}}_{\sf{u}}(k)$-invariant objects. The foliation by the leaves of $V$ will be denoted by $\mathcal{V}$.
A pp-wave is a Lorentzian manifold $(M,V)$ with a parallel null vector field $V$, such that the (integrable) parallel distribution $V^{\perp}$ is flat with respect to the induced connection (this distribution defines a foliation $\mathfrak{F}$ which is totally geodesic). The leaves of $\mathfrak{F}$ are endowed naturally with a $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-structure, which varies smoothly. Indeed, it is well known that the data of a flat affine connection on a smooth manifold is equivalent to have an $(\sf{Aff}(\mathbb{R}^{n+1}),\mathbb{R}^{n+1})$-structure on it. Hence, pp-waves have an induced affine structure on the leaves of $\mathfrak{F}$. However, due to the existence of a parallel vector field $V$ and the fact that $V^{\perp}$ has a natural degenerate Riemannian metric with kernel $\mathbb{R}V$, the leaves have in fact a $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-geometry. So pp-waves give natural examples of $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated manifolds.
2.4. The $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated completeness
This subsection is devoted to the study of the foliated completeness problem, we show the following

### v2

@@ 1.3.2. Affine lightlike geometry
One could relax the $({\sf{L}}_{\sf{u}}(n),\mathbb{R}^{n+1})$-geometry by assuming, instead of an invariant vector field, only an invariant line field. In this case, the structure group is
@@
We refer to this $({\sf{L}}(n),\mathbb{R}^{n+1})$-geometry as the affine lightlike geometry.
Let $(M,l)$ be a ‘weakly pp-wave’ of dimension $n+2$, that is, a Lorentzian manifold with a null parallel line field $l$, for which the (totally geodesic) foliation $\mathfrak{F}$ orthogonal to $l$ is flat with respect to the induced connection. The leaves of $\mathfrak{F}$ have a natural affine lightlike geometry. So weakly pp-waves give natural examples of $(\mathsf{L}(n),\mathbb{R}^{n+1})$-foliated manifolds.
Using essentially the same techniques as in Theorem 1.10, we characterize the completeness in the following way:
@@
Let $(M,\mathfrak{F})$ be a compact $(\mathsf{L}(n),\mathbb{R}^{n+1})$-foliated manifold. The leaves of $\mathfrak{F}$ are $(\mathsf{L}(n),\mathbb{R}^{n+1})$-complete if and only if all geodesics tangent to the $\mathcal{V}$-leaves are affinely (or geodesically) complete, where $\mathcal{V}$ is the $1$-dimensional foliation tangent to the line field induced on any $\mathfrak{F}$-leaf by this structure.
Incompleteness may occur in Theorem 1.13: taking the product of a Hopf circle and a Euclidean torus (Example 3.2) defines such an example.
@@
A natural generalization of the above situation is the following pseudo-Riemannian variant of the $({\sf{L}}_{\sf{u}}(n,0),\mathbb{R}^{n+1})$-geometry. Let $p+q=n$, and define
@@
where $\mathsf{O}(p,q)$ is the orthogonal group of the usual pseudo-Riemannian scalar product on $\mathbb{R}^{p+q}$ of signature $(p,q)$. One can define similarly ${\sf{L}}(p,q)$, the pseudo-Riemannian analogue of ${\sf{L}}(n)$.
Example 1.14 (Higher-index pp-waves).
There are also natural examples of $({\sf{L}}_{\sf{u}}(p,q),\mathbb{R}^{p+q+1})$-foliated manifolds, given by higher-index pp-waves of signature $(p+1,q+1)$. A higher-index pp-wave can be defined similarly to the Lorentzian case as a pseudo-Riemannian manifold of signature $(p+1,q+1)$, admitting a lightlike parallel vector field $V$, such that the distribution $V^{\perp}$ is flat with respect to the induced connection. As in the Lorentzian case, they admit Brinkmann local coordinates (the proof in [27, Paragraph 4.2.4] holds independently of the signature) of the form
$2dudv+H(u,x)du^{2}+\sum_{1}^{p}dx_{i}^{2}-\sum_{1}^{q}dx_{i}^{2}.$
Those spaces possess a natural tangential $({\sf{L}}_{\sf{u}}(p,q),\mathbb{R}^{p+q+1})$-structure along the foliation tangent to $V^{\perp}$. In Lorentzian signature $q+1=1$, the isometries preserving the foliation act as automorphisms of the ${\sf{L}}(p,q)$-structure on the leaves, since in this case $V$ is the unique (up to scale) parallel null vector field tangent to the foliation. This is no longer true in higher index, and this is a fundamental difference between the Lorentzian and higher index pp-waves from the point of view of geometric structures.
Example 1.15 (A compact ${\sf{L}}_{\sf{u}}(1,1)$-foliated manifold).
In dimension $4$, there is a special example of a symmetric pp-wave of index $2$, given by the hyperbolic oscillator group ${\mathsf{Osc}}_{s}:=\mathbb{R}\ltimes{\mathsf{Heis}}_{3}$, where the $\mathbb{R}$-action acts trivially on the center $Z$ of ${\mathsf{Heis}}_{3}$, and by hyperbolic transformations on ${\mathsf{Heis}}_{3}/Z$. The hyperbolic oscillator has a bi-invariant metric $g_{{\mathsf{Osc}}_{s}}$ of signature $(2,2)$. Considering the space $({\mathsf{Osc}}_{s},g_{{\mathsf{Osc}}_{s}})$ with its full isometry group gives what we call the hyperbolic oscillator geometry. The latter geometry admits a unique (up to scale) parallel lightlike vector field $V$, and the orthogonal distribution $V^{\perp}$ is flat. So, by taking a cocompact lattice $\Gamma$ of ${\mathsf{Osc}}_{s}$, one obtains a compact (locally symmetric) ${\sf{L}}_{\sf{u}}(1,1)$-foliated manifold ${\mathsf{Osc}}_{s}/\Gamma$. The $4$-dimensional symmetric pp-waves of index $2$ have been recently explored in [24, 20], where the authors compute the full isometry group and study the existence of compact complete manifolds modeled on them.
The famous Markus conjecture states that an affine structure on a compact manifold is complete as soon as it admits a parallel volume form. The $({\sf{L}}_{\sf{u}}(p,q),\mathbb{R}^{p+q+1})$ geometry and its dual are part of this conjecture, and therefore, compact $({\sf{L}}_{\sf{u}}(p,q),\mathbb{R}^{p+q+1})$-manifolds (and their duals) are conjectured to be complete. However, it turns out that the foliated completeness Problem 1.8 fails in this case. We provide examples of compact $({\sf{L}}_{\sf{u}}(p,q),\mathbb{R}^{p+q+1})$-foliated and $({\sf{L}}_{\sf{u}}^{*}(p,q),\mathbb{R}^{p+q+1})$-foliated manifolds (for any $p,q\geq 1$) with incomplete leaves (Example 4.5 and Example 4.6). The issue arises from the existence of compact manifolds with foliations that admit a tangential flat Lorentzian structure, such that the leaves are incomplete. Or, in other words, from the fact that there is no foliated version of Carrière’s theorem on the completeness of compact flat Lorentzian manifolds, as it is observed in Example 4.4.
For the dual geometry, with $q=1$, we show
@@
Let $M$ be a compact $({\sf{L}}_{\sf{u}}^{*}(n-1,1),\mathbb{R}^{n+1})$-manifold. Then the developing map is injective; in particular, $M$ is Kleinian. If moreover, the fundamental group of $M$ is virtually solvable, then $M$ is complete.
@@
The organization of the paper is different from the order presented in the introduction. Section 2 deals with the $(G,X)$-foliated completeness problem in Item (B), in the case of the affine unimodular lightlike geometry and its dual. We prove Theorem 1.10 and Theorem 1.11. Then we study the case of the affine lightlike geometry in Section 3: we prove Theorem 1.13, and give leafwise-incomplete examples. Section 4 deals with the pseudo-Riemannian variants of the previous structures. In Section 5, we deal with the foliated completeness problem in Item (A), and prove Theorem 5.1. Finally, Section 6 is devoted to the study of the global completeness question in the presence of a null Killing field, focusing on dimension $3$ in Section 7.
Acknowledgment
We would like to warmly thank Abdelghani Zeghib and Souheib Allout for the helpful discussions and valuable comments on this work. The first author is grateful to Ines Kath for her support, encouragement, and for useful discussions and remarks on early drafts of this paper. The second author would like to warmly thank Miguel Sanchez Caja for the helpful discussions. Finally, we thank the referee for the valuable comments and suggestions that helped improve the quality of the presentation.
The second author is supported by the grants, PID2020-116126GB-I00
@@ 2. $(G,X)$-foliated completeness problem: affine unimodular lightlike geometry
Let $G$ be a Lie group, and $X$ a space on which $G$ acts transitively by real analytic diffeomorphisms. Then $X=G/H$, where $H$ is a closed subgroup of $G$. A $(G,X)$-structure on a smooth manifold $M$ is an atlas on $M$, where the chart transitions are restrictions of elements of $G$. A manifold with a $(G,X)$-structure is called a $(G,X)$-manifold. A $(G,X)$-manifold $M$ gives rise to a local diffeomorphism ${\mathsf{Dev}}$, called the developing map, from the universal cover to the model space $X$, along with a representation $\rho:\pi_{1}(M)\to G$, called the holonomy representation (see [14, p. 100] or [34, p. 34]). The developing map and the holonomy representation satisfy the following equivariance property
${\mathsf{Dev}}(\gamma x)=\rho(\gamma){\mathsf{Dev}}(x),$
@@
Let $M$ be a $(G,X)$-manifold. We say that $M$ is complete if the developing map ${\mathsf{Dev}}:\widetilde{M}\to X$ is a covering map.
@@ 2.2. Tangential $(G,X)$-foliation
We introduce here the notion of a $(G,X)$-foliated manifold. A smooth manifold $M$ with a foliation $\mathfrak{F}$ is referred to as a foliated manifold, denoted by the pair $(M,\mathfrak{F})$.
Definition 2.3 (Tangential $(G,X)$-foliation).
@@
We say that a manifold $M$ has a tangential $(G,X)$-foliation if $M$ is a foliated manifold such that the leaves of $\mathfrak{F}$ are modeled on $(G,X)$. More precisely, for every $p\in M$, there exists an open neighborhood $U$ and a diffeomorphism (into the image) $\varphi:U\to\mathbb{R}^{k}\times X$ such that every leaf of the foliation of $U$ (induced by $\mathfrak{F}$) is mapped to a level $(t,V_{t})\subset\mathbb{R}^{k}\times X$. And if $(\varphi_{i},U_{i})$ and $(\varphi_{j},U_{j})$ are two such maps, the transition map $\varphi_{ij}:=\varphi_{j}\varphi_{i}^{-1}$ has the following form $(\psi(t),g_{t}(x))$ where $g_{t}\in G$ and $x\in V_{t}$ (compare [11, 19, 25]).
Convention: We call a $(G,X)$-foliated manifold a manifold $M$ having a tangential $(G,X)$-foliation, as defined above.
In addition to the examples provided in the introduction, we present here some natural examples of foliations with a tangential structure:
Example 2.4 (Vector fields).
The data of a non-vanishing vector field on $M$ is equivalent to giving a $1$-dimensional foliation with a tangential $(\mathbb{R},\mathbb{R})$-structure. The structure on the leaves is complete if and only if the vector field is complete. This is the case when $M$ is compact.
Example 2.5 (Lightlike foliations of Lorentzian surfaces).
An orientable Lorentzian surface has two totally geodesic lightlike foliations. Each foliation naturally carries a tangential affine structure, modeled on $(\sf{Aff}(\mathbb{R}),\mathbb{R})$. The local affine parameter is given by the geodesic parameter. The completeness of this affine structure is equivalent to the completeness of the lightlike geodesics tangent to the foliation.
Example 2.6 (A tangential Minkowski structure).
Consider $\mathbb{R}^{2}$ with the action of the hyperbolic matrix $\varphi:=\begin{pmatrix}2&1\\ 1&1\end{pmatrix}$ (this means that it has two distinct real eigenvalues). There is a Lorentzian flat metric on $\mathbb{R}^{2}$ preserved by $\varphi$ (take two $1$-forms $du$ and $dv$ defining the two eigendirections of $\varphi$, and set $g:=dudv$). Define the flat $2$-torus $\mathbb{T}^{2}:=\mathbb{R}^{2}/\mathbb{Z}^{2}$, with $\overline{\varphi}$ the induced action of $\varphi$. The mapping torus $\mathbb{T}_{\varphi}:=\mathbb{T}^{2}\times[0,1]/(x,0)\sim(\overline{\varphi}(x),1)$ has a natural foliation with a tangential Minkowski structure, whose leaves are given by the $\mathbb{T}^{2}$-fibers.
Example 2.7 (A tangential hyperbolic structure).
Consider the group ${\mathsf{Sol}}\cong\mathbb{R}\ltimes\mathbb{R}^{2}$, with $\mathbb{R}$ acting on $\mathbb{R}^{2}$ by linear hyperbolic transformations $\begin{pmatrix}e^{t}&0\\ 0&e^{-t}\end{pmatrix}.$ We have a short exact sequence $\sf{Aff}(\mathbb{R})\to{\mathsf{Sol}}\to\mathbb{R}$. The left action of $\sf{Aff}(\mathbb{R})$ on ${\mathsf{Sol}}$ defines a foliation $\mathfrak{F}$ with a tangential $(\sf{Aff}(\mathbb{R}),\sf{Aff}(\mathbb{R}))$-structure. On the other hand, $\sf{Aff}(\mathbb{R})$ acts simply transitively on the hyperbolic plane $\mathbb{H}^{2}$, inducing a natural hyperbolic structure on the leaves of $\mathfrak{F}$. The foliation is right invariant, so by taking a cocompact lattice $\Gamma$ of ${\mathsf{Sol}}$, we obtain a compact manifold ${\mathsf{Sol}}/\Gamma$ with a tangential hyperbolic foliation.
Example 2.8.
(Tangential $(G,G)$-foliations) This is defined similarly to the previous example. Let $L$ be a connected Lie group, and $G$ a connected closed subgroup of $L$. Then $L$ is naturally a $(G,G)$-foliated manifold. We obtain a compact $(G,G)$-foliated manifold if $L$ has a cocompact lattice.
@@
Pullback of $G$-invariant tensor fields. Consider $(M,\mathfrak{F})$ a $(G,X)$-foliated manifold. Let $T$ be a tensor field on $X$ which is $G$-invariant. One can associate a tensor field $T_{M}$ on $M$ in the following way. For any point $p\in M$, we take a chart around $p$ and we pullback the tensor field $T$ in a neighborhood of $p$. Since the transition maps, when restricted to a leaf, are restrictions of elements of $G$, and $T$ is $G$-invariant, we get a well defined tensor field $T_{M}$ on $M$. Moreover, by construction, $T_{M}$ is continuous on $M$. We will refer to $T_{M}$ as the “pullback” of $T$.
Remark 2.9 (Leafwise completeness: tautological cases).
There are two situations, one geometric and the other topological, in which any compact $(G,X)$-foliated manifold $(M,\mathfrak{F})$ is leafwise $(G,X)$-complete. The first situation is when the stabilizer in $G$ of some point in $X$ is compact; equivalently, $X$ admits a $G$-invariant Riemannian metric. In this case, taking the pullback of this metric defines a Riemannian metric on $T\mathfrak{F}$. Then, the completeness follows from Fact 2.15. The other situation is when all the leaves are simply connected, or more generally, when the holonomy representation of any leaf is trivial. Then $T\mathfrak{F}$ also supports a Riemannian metric, since the transition maps of the geometric structure are just restrictions of the identity map of $G$, and the leaves are therefore $(G,X)$-complete.
@@
Recall that the affine unimodular lightlike geometry is the one modeled on $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$, where $\mathsf{L}_{\sf{u}}(n)$ is the subgroup of $\sf{Aff}(\mathbb{R}^{n+1})$ preserving the degenerate Riemannian metric $h:=dx_{2}^{2}+dx_{3}^{2}+\ldots+dx_{n+1}^{2}$  and the null vector field $V:=\partial_{x_{1}}$. In this subsection, $V$ and $h$ will always refer to these ${\sf{L}}_{\sf{u}}(n)$-invariant objects. The foliation by the leaves of $V$ will be denoted by $\mathcal{V}$.
2.4. $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated completeness
This subsection is devoted to the study of the foliated completeness problem. We show the following